# Zero-shot Generalization in Inventory Management: Train, then Estimate and Decide

### v2

@@ 4 Problem Description
Assumption 1 enables the decomposition of the problem into a collection of independent and stationary MDPs, each characterized by a distinct parameterization. [^1] Even if a parameterization induces nonstationary dynamics, it can still define a stationary (time-homogeneous) MDP when the nonstationarity is structured and incorporated into the state space (Kallenberg,, 2011, Sec. 2.3). This approach is common in inventory management, where optimization strategies are typically based on current estimates of problem parameters (Axsäter,, 2015). In fact, it is akin to a rolling horizon strategy: the parameterization is assumed to be stationary for the time being, while at each time step, changes in parameters are monitored, and values are potentially re-estimated if they are unknown.
@@
This assumption is reasonable because the operational environment is typically well understood at a higher level, parameters are often naturally bounded (e.g., non-negativity of demand), and such assumptions are commonly employed in the online learning literature (see Lyu et al.,, 2024, and the references therein). We later define and utilize a probable parameterization space and establish a parameter sampling function to conceptualize the Super-MDP.
@@ 5 Train, then Estimate and Decide
The main novelty of this approach is that the training step, during which the policy is optimized, precedes the estimation step. This contrasts with current prevalent approaches, which first estimate parameters and then optimize based on parameter estimates; see Agrawal and Jia, (2022); Lyu et al., (2024); Chen et al., 2023a for similar applications covering online learning in inventory management. We provide detailed descriptions of the different phases in the subsequent sections.
@@ 5.1 Train
The construction of the probable parameter space $\hat{\mathcal{P}}$ and the probability distribution $\hat{\mathcal{H}}$ defined over it may influence the generalization performance of the GCA. By defining $\hat{\mathcal{P}}$ as a superset of the true parameter set $\mathcal{P}$, reasonably similar parameterizations are guaranteed to be available for sampling during training. This allows the agent to encounter a broad spectrum of representative problem instances. The rationale becomes particularly evident under the Lipschitz conditions for Super-MDPs: when these are satisfied, small perturbations in parameterizations induce proportionally small changes in transition dynamics and costs; thus, controlling the effect of parameter variation.
@@
where $\|p-q\|_{1}:=\sum_{x\in\mathcal{S}}|p(x)-q(x)|$ is the $\ell_{1}$ distance between the two transition probability distributions over next states for the same $(\mathbf{s},a)$.
@@
The Lipschitz property provides actionable guarantees for the ZSG capabilities of a GCA. Specifically, the Lipschitz conditions impose a smoothness constraint on how transition probabilities and reward functions vary with the problem parameters. This smoothness supports predictable and continuous system behavior as parameters change. When these conditions hold, an agent that learns a decision-making policy from collected problem instances within a parameter space is better positioned to anticipate and perform well for closely related yet unseen instances (we provide a quantitative analysis in §5.2 when the true parameters are unknown and should be estimated). In this case, extrapolation of learned behaviors to new instances is facilitated by the Lipschitz guarantee that dynamics and rewards will not differ drastically, thereby reducing the need for additional training. Consequently, a policy effective in one instance is more likely to succeed in similar ones, strengthening the GCA’s performance in a ZSG setting (see Xu and Mannor,, 2012, for a related discussion).
@@
It is important to emphasize, however, that the Lipschitz property alone does not guarantee ZSG, and ZSG may still be observed empirically for non-Lipschitz Super-MDPs, but in such cases, no formal guarantees can be established, such as the one we will derive in §5.2. The realized test-time performance gap also depends on (i) the extent to which $\hat{\mathcal{P}}$ is adequately covered by training samples from $\hat{\mathcal{H}}$, (ii) the accuracy of parameter estimation at deployment, and (iii) the degree of optimization or approximation error in the learned policy. Poor coverage (e.g., a narrowly concentrated $\hat{\mathcal{H}}$), inaccurate parameter estimates, or insufficient training can all degrade performance even if the underlying Super-MDP is Lipschitz.
@@
We now explore algorithmic strategies for training GCAs under ZSG and utilize DRL as our primary methodological tool due to its proven success in handling large-scale sequential decision-making problems. According to Kirk et al., (2023), supported by numerous references in their paper, methods addressing ZSG challenges in DRL can be categorized into three main areas: DRL-specific improvements, enhancing similarity between training and testing data, and addressing differences between training and testing environments. We begin by examining DRL-specific strategies for algorithm adoption.
The DRL domain has advanced significantly since the seminal introduction of the Deep Q-Network (DQN) algorithm by Mnih et al., (2013). However, designing algorithms with ZSG capability remains challenging, as highlighted by Kirk et al., (2023). They identify several promising directions to enhance ZSG, including: (i) iterative policy distillation into reinitialized networks to mitigate non-stationarity (i.e., the shifting data distribution caused by the evolving policy during training), (ii) architectural separation of policy and value networks for more targeted optimization, and (iii) model-based approaches to reduce sample complexity and improve generalization.
Given these considerations, the Deep Controlled Learning (DCL) algorithm (Temizöz et al.,, 2025) emerges as a natural candidate for our application. DCL is specifically designed for environments influenced by exogenous stochastic factors, such as demand and lead time distributions, making it well suited for solving MDPs whose parameterizations encode stochastic processes (cf. §3). From a methodological perspective, DCL fits the conditions for ZSG capability: it is a model-based approximate policy iteration algorithm that iteratively refines policies by distilling them into reinitialized policy networks and avoiding the need for a value network during optimization through efficient simulations.
Since DCL does not inherently support the Super-MDP formulation for training GCAs under ZSG, we extend it by adopting the strategies outlined by Kirk et al., (2023): enhancing similarity between training and testing data, and addressing differences between training and testing environments. Algorithm 1 details this adaptation, referred to as Deep Controlled Learning for Super-MDPs (Super-DCL).
@@
12:     end for
13:   end for
14:  end for
15:  $\pi_{i+1}=Classifier(N_{\theta},\mathcal{K}_{i})$
16:end for
17:Output: $\pi_{1},\dots,\pi_{\bar{n}}$
@@
Before presenting the details of this adaptation, we briefly summarize the algorithm’s core mechanics. Super-DCL improves an initial policy (denoted $\pi_{S,0}$ or simply $\pi_{0}$) through successive iterations of approximate policy improvement, framing reinforcement learning as a classification problem. At each iteration $i$, the algorithm builds a dataset $\mathcal{K}_{i}$ of parameterization–state pairs together with their estimated optimal actions. To construct $\mathcal{K}_{i}$, parameterizations are sampled from the probable parameter generator $\mathbf{p}\sim\hat{\mathcal{H}}$, and each sampled $\mathbf{p}$ induces an MDP $\mathcal{M}^{\mathbf{p}}=\mathcal{F}(\mathbf{p})$ (Line 6). From this MDP, states are collected and a simulator estimates the optimal action $\hat{\pi}_{i}^{+}$ by evaluating candidate actions through $M$ rollouts of depth $H$ (Line 9), thereby providing the reinforcement signal that drives policy improvement. Each labeled pair $((\mathbf{s},\mathbf{p}),\hat{\pi}_{i}^{+}(\mathbf{s},\mathbf{p}))$ is then added to $\mathcal{K}_{i}$ (Line 10). Once dataset $\mathcal{K}_{i}$ is complete, the policy is updated by training a neural network classifier from scratch, yielding the new policy $\pi_{i+1}$ (Line 15); the network learns to map each parameterization–state pair $(\mathbf{s},\mathbf{p})$ to its estimated optimal action $\hat{\pi}_{i}^{+}(\mathbf{s},\mathbf{p})$. Repeating this process for $\bar{n}$ iterations yields a sequence of increasingly refined policies $\pi_{1},\dots,\pi_{\bar{n}}$, with the final output $\pi_{\bar{n}}$ serving as an approximation of the optimal policy $\hat{\pi}_{S}$.
@@
Kirk et al., (2023) emphasize that increasing the similarity between training and testing data can substantially reduce the generalization gap and improve performance during testing. This similarity can be achieved by ensuring that the testing environment lies within the range of environments encountered during training. Such alignment can be facilitated through two complementary strategies: extensive sampling, where we utilize a large dataset during training, and domain randomization, which for the DCL algorithm can be supported by (i) the introduction of a new hyperparameter $R$ that triggers resampling of new parameterizations after $R$ samples have been collected for a specific parameterization, and (ii) the design of the probable parameter generator $\hat{\mathcal{H}}$ to provide uniform coverage of the probable space $\hat{\mathcal{P}}$, thereby exposing the agent to a wide variety of environments. Ensuring such uniformity is important: it increases the likelihood that unseen test instances fall close to trained ones, thereby strengthening ZSG. The potential benefits of these strategies are illustrated in Figure 2, which shows how broad sampling and randomized parameterization design expand the coverage of the true, unknown parameter space and thus enhance ZSG performance.
@@
Furthermore, ZSG can be improved by training the agent to exploit differences among the environments it encounters. By incorporating parameters as features during training (Line 10), the GCA learns to identify and adapt to the dynamics of each task ($\mathcal{M}^{\mathbf{p}}$). This strategy allows the agent to effectively apply these learned distinctions during testing, thereby bridging the gap between the training and testing environments (Kirk et al.,, 2023). Such awareness during the learning process serves as an inductive bias, enabling the agent to accurately identify and respond to environmental characteristics at test time.
@@
The main novelty in Super-DCL lies in moving beyond training an MDP for a fixed parameterization $\mathbf{p}^{j}$ to varying the parameterization using a previously defined sampling function $\hat{H}$ and reinitializing the system after $R$ sampling periods with the fixed parameterization $\mathbf{p}^{j}$ (Lines 5-11). Moreover, the policy $\pi$ not only takes states as input but also incorporates the parameterization used to generate the training samples (Line 10). For a comprehensive explanation of DCL, including the details of the sampling process, parallelization among threads (Line 4), the warm-up step for sampling a state (Line 8), and the $Simulator$ (Line 9) and $Classifier$ (Line 15) algorithms, we direct readers to Temizöz et al., (2025).
@@ 5.2 Estimate
An estimator $\mathcal{Y}:\mathcal{O}_{t}\to\hat{\mathcal{P}}$ is consistent if
$d_{P}(\mathbf{p},\hat{\mathbf{p}}_{t})\ \xrightarrow[t\to\infty]{\mathbb{P}}\ 0,\quad\text{that is,}\quad\forall\varepsilon>0:\ \lim_{t\to\infty}\Pr\big{(}d_{P}(\mathbf{p},\hat{\mathbf{p}}_{t})>\varepsilon\big{)}=0.$
@@
This proof follows from the Simulation Lemma (Kearns and Singh,, 2002) applied to the Super-MDPs; and can be found in Appendix A. ∎
@@ 6 Periodic Review Inventory Control
In general, for problems in this class, the sequence of events for any given task and time point $t$ unfolds as follows. Let $OH_{t}$ denote the on-hand inventory at the beginning of the period $t$. At the start of period $t$, a new shipment of inventory $q_{t}$ is delivered and added to $OH_{t}$. Then, the DM reviews the inventory and places a new order $a_{t}$. If the supplier allows for instant delivery, a portion of this order, $q_{t}^{0}$, may be immediately added to the on-hand inventory. The rest of the orders will be received according to a lead time distribution. The random demand $D_{t}$ is realized, and the DM observes the sales, which are the minimum of $D_{t}$ and the available inventory $OH_{t}+q_{t}+q_{t}^{0}$. The unmet demands are lost. At the end of period $t$, the DM incurs costs based on the unsatisfied demands and the on-hand inventory they have. In particular, the cost at the end of the period is: $C_{t}=h(OH_{t}+q_{t}+q_{t}^{0}-D_{t})^{+}+p(D_{t}-(OH_{t}+q_{t}+q_{t}^{0}))^{+}$, where $h$ is the holding cost of an item in inventory, and $p$ is the penalty cost for each unmet demand. However, demand may be censored in the event of stockouts, where only sales data is accessible, not the actual demand. In this case, penalty cost is not observable, and Lyu et al., (2024) instead adopt maximizing the profit ($G_{t}$) as an objective, where $p$ is interpreted as per-unit sales revenue. Adapting this objective to our setting, we write: $G_{t}=p\min(D_{t},OH_{t}+q_{t}+q_{t}^{0})-h(OH_{t}+q_{t}+q_{t}^{0}-D_{t})^{+}$, which is applicable when the demand distribution is unknown and censored.
@@
We next briefly explain the details of the demand and supply processes that determine the amounts $D_{t}$, $q_{t}$ and $q_{t}^{0}$. The demand $D_{t}$ at each time step $t$ is considered a discrete random variable within the bounds of $0,D_{\max}]$ and follows a structured pattern. We model this demand as occurring in cycles of constant length $K$, where $K$ may span from $1$ to $K_{max}$. The demands are independent across the periods but not identically distributed within a cycle, allowing for a variety of distribution patterns. Mathematically, the demands $D_{t}$ can be grouped into $K$ subsets according to the remainder when $t$ is divided by $K$ (i.e., $t\mod K$), with each subset of demands being generated independently from its specific distribution (see Gong and Simchi-Levi,, 2024, for a similar demand process analysis).
The supply process in our inventory system, determining the amounts $q_{t}$ and $q_{t}^{0}$, is influenced by stochastic lead times. To model this process, we reference two prevalent approaches in the literature: orders can cross, where an order that is placed after another order may be received before that order (order crossover, see Stolyar and Wang,, 2022), and orders cannot cross (see Kaplan, (1970)). In the former case, the lead times are i.i.d., while in the latter case orders are received in the same sequence in which they were placed. We incorporate both cases as they have practical relevance; see Andaz et al., (2023) for order crossing, and Bai et al., 2023a for the latter case and a general analysis on stochastic lead times.
@@
Our problem class includes instances with a wide range of cost parameters, demand distributions (including cyclic demand), and lead time distributions (with or without order crossover). When constructing a Super-MDP for this problem during the Train phase, we focus on strategies to maintain the Lipschitz conditions of the Super-MDP (see Definition 4). The probable parameter space and the associated parameter generation function are created accordingly. We define bounds for the cost parameter $p$ and parameters related to demand and lead time distributions, and we let demand and lead time distributions arise from two-moment fits (following Adan et al.,, 1995) and other techniques. For details, we refer to Appendix C, which contains a detailed description of the construction of the Super-MDP as well as the subsequent training of the GCA.
In the Estimate phase, Theorem 1 encourages the use of any effective technique to accurately estimate the true parameters. We assume that the holding cost $h$ and penalty cost $p$ are directly observable, while the demand distribution $\boldsymbol{\zeta}$ and lead time distribution $\boldsymbol{\tau}$ are unknown and should be estimated from historical data. When demands are censored, where only sales data is available, true demand can be systematically underestimated, potentially causing continuous understocking (Huh and Rusmevichientong,, 2009). To mitigate this, we apply the Kaplan–Meier estimator (Kaplan and Meier,, 1958) to estimate the demand distribution under censored observations; similar adaptations appear in Huh et al., (2011) and Lyu et al., (2024).

### v3

@@ 4 Problem Description
Assumption 1 enables the decomposition of the problem into a collection of independent and stationary MDPs, each characterized by a distinct parameterization. [^2] Even if a parameterization induces nonstationary dynamics, it can still define a stationary (time-homogeneous) MDP when the nonstationarity is structured and incorporated into the state space (Kallenberg, 2011, Sec. 2.3). This approach is common in inventory management, where optimization strategies are typically based on current estimates of problem parameters (Axsäter, 2015). In fact, it is akin to a rolling horizon strategy: the parameterization is assumed to be stationary for the time being, while at each time step, changes in parameters are monitored, and values are potentially re-estimated if they are unknown.
@@
This assumption is reasonable because the operational environment is typically well understood at a higher level, parameters are often naturally bounded (e.g., non-negativity of demand), and such assumptions are commonly employed in the online learning literature (see Lyu et al., 2024, and the references therein). We later define and utilize a probable parameterization space and establish a parameter sampling function to conceptualize the Super-MDP.
@@ 5 Train, then Estimate and Decide
The main novelty of this approach is that the training step, during which the policy is optimized, precedes the estimation step. This contrasts with current prevalent approaches, which first estimate parameters and then optimize based on parameter estimates; see Agrawal and Jia (2022); Lyu et al. (2024); Chen et al. (2023a) for similar applications covering online learning in inventory management. We provide detailed descriptions of the different phases in the subsequent sections.
@@ 5.1 Train
The construction of the probable parameter space $\hat{\mathcal{P}}$ and the probability distribution $\hat{\mathcal{H}}$ defined over it may influence the generalization performance of the GCA. By defining $\hat{\mathcal{P}}$ as a superset of the true parameter set $\mathcal{P}$, reasonably similar parameterizations are guaranteed to be available for sampling during training. This allows the agent to encounter a broad spectrum of representative problem instances. The rationale becomes particularly evident under the Lipschitz conditions for Super-MDPs: when these are satisfied, small perturbations in parameterization induce proportionally small changes in transition dynamics and costs; thus, controlling the effect of parameter variation.
@@
where $\|p-q\|_{1}:=\sum_{x\in\mathcal{S}}|p(x)-q(x)|$ is the $\ell_{1}$ distance between the two transition probability distributions over the next states for the same $(\mathbf{s},a)$.
@@
The Lipschitz property provides actionable guarantees for the ZSG capabilities of a GCA. Specifically, the Lipschitz conditions impose a smoothness constraint on how transition probabilities and reward functions vary with the problem parameters. This smoothness supports predictable and continuous system behavior as parameters change. When these conditions hold, an agent that learns a decision-making policy from collected problem instances within a parameter space is better positioned to anticipate and perform well for closely related yet unseen instances (we provide a quantitative analysis in §5.2 when the true parameters are unknown and should be estimated). In this case, extrapolation of learned behaviors to new instances is facilitated by the Lipschitz guarantee that dynamics and rewards will not differ drastically, thereby reducing the need for additional training. Consequently, a policy effective in one instance is more likely to succeed in similar ones, strengthening the GCA’s performance in a ZSG setting (see Xu and Mannor, 2012, for a related discussion).
@@
It is important to emphasize, however, that the Lipschitz property alone does not guarantee ZSG, and ZSG may still be observed empirically for non-Lipschitz Super-MDPs, but in such cases, no formal guarantees can be established, such as the one we derive in §5.2. The realized test-time performance gap also depends on (i) the extent to which $\hat{\mathcal{P}}$ is adequately covered by training samples from $\hat{\mathcal{H}}$, (ii) the accuracy of parameter estimation at deployment, and (iii) the degree of optimization or approximation error in the learned policy. Poor coverage (e.g., a narrowly concentrated $\hat{\mathcal{H}}$), inaccurate parameter estimates, or insufficient training can all degrade performance even if the underlying Super-MDP is Lipschitz.
@@
We now explore algorithmic strategies for training GCAs under ZSG and utilize DRL as our primary methodological tool due to its proven success in handling large-scale sequential decision-making problems. According to Kirk et al. (2023), supported by numerous references in their paper, methods addressing ZSG challenges in DRL can be categorized into three main areas: DRL-specific improvements, enhancing similarity between training and testing data, and addressing differences between training and testing environments. We begin by examining DRL-specific strategies for algorithm adoption.
The DRL domain has advanced significantly since the seminal introduction of the Deep Q-Network (DQN) algorithm by Mnih et al. (2013). However, designing algorithms with ZSG capability remains challenging, as highlighted by Kirk et al. (2023). They identify several promising directions to enhance ZSG, including: (i) iterative policy distillation into reinitialized networks to mitigate non-stationarity (i.e., the shifting data distribution caused by the evolving policy during training), (ii) architectural separation of policy and value networks for more targeted optimization, and (iii) model-based approaches to reduce sample complexity and improve generalization.
Given these considerations, the Deep Controlled Learning (DCL) algorithm (Temizöz et al., 2025) emerges as a natural candidate for our application. DCL is specifically designed for environments influenced by exogenous stochastic factors, such as demand and lead time distributions, making it well suited for solving MDPs whose parameterizations encode stochastic processes (cf. §3). From a methodological perspective, DCL fits the conditions for ZSG capability: it is a model-based approximate policy iteration algorithm that iteratively refines policies by distilling them into reinitialized policy networks and avoiding the need for a value network during optimization through efficient simulations.
Since DCL does not inherently support the Super-MDP formulation for training GCAs under ZSG, we extend it by adopting the strategies outlined by Kirk et al. (2023): enhancing similarity between training and testing data, and addressing differences between training and testing environments. Algorithm 1 details this adaptation, referred to as Deep Controlled Learning for Super-MDPs (Super-DCL).
@@
12:  $\pi_{i+1}=Classifier(N_{\theta},\mathcal{K}_{i})$
13:Output: $\pi_{1},\dots,\pi_{\bar{n}}$
@@
Before presenting the details of this adaptation, we briefly summarize the algorithm’s core mechanics. Super-DCL improves an initial policy (denoted $\pi_{S,0}$ or simply $\pi_{0}$) through successive iterations of approximate policy improvement, framing reinforcement learning as a classification problem. At each iteration $i$, the algorithm builds a dataset $\mathcal{K}_{i}$ of parameterization–state pairs together with their estimated optimal actions. To construct $\mathcal{K}_{i}$, parameterizations are sampled from the probable parameter generator $\mathbf{p}\sim\hat{\mathcal{H}}$, and each sampled $\mathbf{p}$ induces an MDP $\mathcal{M}^{\mathbf{p}}=\mathcal{F}(\mathbf{p})$ (Line 6). From this MDP, states are collected and a simulator estimates the optimal action $\hat{\pi}_{i}^{+}$ by evaluating candidate actions through $M$ rollouts of depth $H$ (Line 9), thereby providing the reinforcement signal that drives policy improvement. Each labeled pair $((\mathbf{s},\mathbf{p}),\hat{\pi}_{i}^{+}(\mathbf{s},\mathbf{p}))$ is then added to $\mathcal{K}_{i}$ (Line 10). Once dataset $\mathcal{K}_{i}$ is complete, the policy is updated by training a neural network classifier from scratch, yielding the new policy $\pi_{i+1}$ (Line 12); the network learns to map each parameterization–state pair $(\mathbf{s},\mathbf{p})$ to its estimated optimal action $\hat{\pi}_{i}^{+}(\mathbf{s},\mathbf{p})$. Repeating this process for $\bar{n}$ iterations yields a sequence of increasingly refined policies $\pi_{1},\dots,\pi_{\bar{n}}$, with the final output $\pi_{\bar{n}}$ serving as an approximation of the optimal policy $\hat{\pi}_{S}$.
@@
Kirk et al. (2023) emphasize that increasing the similarity between training and testing data can substantially reduce the generalization gap and improve performance during testing. This similarity can be achieved by ensuring that the testing environment lies within the range of environments encountered during training. Such alignment can be facilitated through two complementary strategies: extensive sampling, where we utilize a large dataset during training, and domain randomization, which for the DCL algorithm can be supported by (i) the introduction of a new hyperparameter $R$ that triggers resampling of new parameterizations after $R$ samples have been collected for a specific parameterization, and (ii) the design of the probable parameter generator $\hat{\mathcal{H}}$ to provide uniform coverage of the probable space $\hat{\mathcal{P}}$, thereby exposing the agent to a wide variety of environments. Ensuring such uniformity is important: it increases the likelihood that unseen test instances fall close to trained ones, thereby strengthening ZSG. The potential benefits of these strategies are illustrated in Figure 2, which shows how broad sampling and randomized parameterization design expand the coverage of the true, unknown parameter space and thus enhance ZSG performance.
@@
Furthermore, ZSG can be improved by training the agent to exploit differences among the environments it encounters. By incorporating parameters as features during training (Line 10), the GCA learns to identify and adapt to the dynamics of each task ($\mathcal{M}^{\mathbf{p}}$). This strategy allows the agent to effectively apply these learned distinctions during testing, thereby bridging the gap between the training and testing environments (Kirk et al., 2023). Such awareness during the learning process serves as an inductive bias, enabling the agent to accurately identify and respond to environmental characteristics at test time.
@@
The main novelty in Super-DCL lies in moving beyond training an MDP for a fixed parameterization $\mathbf{p}^{j}$ to varying the parameterization using a previously defined sampling function $\hat{H}$ and reinitializing the system after $R$ sampling periods with the fixed parameterization $\mathbf{p}^{j}$ (Lines 5-11). Moreover, the policy $\pi$ not only takes states as input but also incorporates the parameterization used to generate the training samples (Line 10). For a comprehensive explanation of DCL, including the details of the sampling process, parallelization among threads (Line 4), the warm-up step for sampling a state (Line 8), and the $Simulator$ (Line 9) and $Classifier$ (Line 12) algorithms, we direct readers to Temizöz et al. (2025).
@@ 5.2 Estimate
An estimator $\mathcal{Y}:\mathcal{O}_{t}\to\hat{\mathcal{P}}$ is consistent if $d_{P}(\mathbf{p},\hat{\mathbf{p}}_{t})\ \xrightarrow[t\to\infty]{\mathbb{P}}\ 0,\quad\text{that is,}\quad\forall\varepsilon>0:\ \lim_{t\to\infty}\Pr\big(d_{P}(\mathbf{p},\hat{\mathbf{p}}_{t})>\varepsilon\big)=0$.
@@
This proof follows from the Simulation Lemma (Kearns and Singh, 2002) applied to the Super-MDPs; and can be found in Appendix A. ∎
@@ 6 Periodic Review Inventory Control
In general, for problems in this class, the sequence of events for any given task and time point $t$ unfolds as follows. Let $OH_{t}$ denote the on-hand inventory at the beginning of the period $t$. At the start of period $t$, a new shipment of inventory $q_{t}$ is delivered and added to $OH_{t}$. Then, the DM reviews the inventory and places a new order $a_{t}$. If the supplier allows for instant delivery, a portion of this order, $q_{t}^{0}$, may be immediately added to the on-hand inventory. The rest of the orders will be received according to a lead time distribution. The random demand $D_{t}$ is realized, and the DM observes the sales, which are the minimum of $D_{t}$ and the available inventory $OH_{t}+q_{t}+q_{t}^{0}$. The unmet demands are lost. At the end of period $t$, the DM incurs costs based on the unsatisfied demands and the on-hand inventory they have. In particular, the cost at the end of the period is: $C_{t}=h(OH_{t}+q_{t}+q_{t}^{0}-D_{t})^{+}+p(D_{t}-(OH_{t}+q_{t}+q_{t}^{0}))^{+}$, where $h$ is the holding cost of an item in inventory, and $p$ is the penalty cost for each unmet demand. However, demand may be censored in the event of stockouts, where only sales data is accessible, not the actual demand. In this case, penalty cost is not observable, and Lyu et al. (2024) instead adopt maximizing the profit ($G_{t}$) as an objective, where $p$ is interpreted as per-unit sales revenue. Adapting this objective to our setting, we write: $G_{t}=p\min(D_{t},OH_{t}+q_{t}+q_{t}^{0})-h(OH_{t}+q_{t}+q_{t}^{0}-D_{t})^{+}$, which is applicable when the demand distribution is unknown and censored.
@@
We next briefly explain the details of the demand and supply processes that determine the amounts $D_{t}$, $q_{t}$ and $q_{t}^{0}$. The demand $D_{t}$ at each time step $t$ is considered a discrete random variable within the bounds of $0,D_{\max}]$ and follows a structured pattern. We model this demand as occurring in cycles of constant length $K$, where $K$ may span from $1$ to $K_{max}$. The demands are independent across the periods but not identically distributed within a cycle, allowing for a variety of distribution patterns. Mathematically, the demands $D_{t}$ can be grouped into $K$ subsets according to the remainder when $t$ is divided by $K$ (i.e., $t\mod K$), with each subset of demands being generated independently from its specific distribution (see Gong and Simchi-Levi, 2024, for a similar demand process analysis).
The supply process in our inventory system, determining the amounts $q_{t}$ and $q_{t}^{0}$, is influenced by stochastic lead times. To model this process, we reference two prevalent approaches in the literature: orders can cross, where an order that is placed after another order may be received before that order (order crossover, see Stolyar and Wang, 2022), and orders cannot cross (see Kaplan (1970)). In the former case, the lead times are i.i.d., while in the latter case orders are received in the same sequence in which they were placed. We incorporate both cases as they have practical relevance; see Andaz et al. (2023) for order crossing, and Bai et al. (2023) for the latter case and a general analysis on stochastic lead times.
@@
Our problem class includes instances with a wide range of cost parameters, demand distributions (including cyclic demand), and lead time distributions (with or without order crossover). When constructing a Super-MDP for this problem during the Train phase, we focus on strategies to maintain the Lipschitz conditions of the Super-MDP (see Definition 4). The probable parameter space and the associated parameter generation function are created accordingly. We define bounds for the cost parameter $p$ and parameters related to demand and lead time distributions, and we let demand and lead time distributions arise from two-moment fits (following Adan et al., 1995) and other techniques. For details, we refer to Appendix C, which contains a detailed description of the construction of the Super-MDP as well as the subsequent training of the GCA.
In the Estimate phase, Theorem 1 encourages the use of any effective technique to accurately estimate the true parameters. We assume that the holding cost $h$ and penalty cost $p$ are directly observable, while the demand distribution $\boldsymbol{\zeta}$ and lead time distribution $\boldsymbol{\tau}$ are unknown and should be estimated from historical data. When demands are censored, where only sales data is available, true demand can be systematically underestimated, potentially causing continuous understocking (Huh and Rusmevichientong, 2009). To mitigate this, we apply the Kaplan–Meier estimator (Kaplan and Meier, 1958) to estimate the demand distribution under censored observations; similar adaptations appear in Huh et al. (2011) and Lyu et al. (2024).